OpenStudy (anonymous):

Using the substitution u = 1+(x)^1/2 find:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ 1+\sqrt{x} }dx\]

hartnn (hartnn):

if u= 1+x^(1/2) what is 'x' in terms of u ?

OpenStudy (raden):

u = 1 + sqrt(x) ----> sqrt(x) = u - 1 du = 1/2sqrt(x) dx 2 sqrt(x) du = dx 2(u-1) du = dx now, ur integral can be :